OpenStudy (anonymous):

Clara ate 2/5 of an orange. What percent of the orange did Clara eat?

OpenStudy (anonymous):

40 percent

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 } \approx \frac{ 4 }{ 10 } \approx 40percent\]

OpenStudy (anonymous):

40

OpenStudy (mathstudent55):

2 is to 5 as x is to 100% \( \dfrac{2}{5} = \dfrac{x}{100} \) Cross multiply: 5x = 200 x = 40 Answer: 40%